Complete the square and find the indefinite integral (Note: using Trigonometric Substitution);
Answer #1

aa-at-1-aa41] +1 - [2.1] 4 22-2?= 1-6-1) das pet nata sint ax= sintal dx = cost at costat [ {int +1) I de sintet inte taste)

sint = 100s (24) we know that **dx = $(n-cos@y) sint +1) at Vanza JC4-couet) 4 2.kint si at -18-Coulet)4 sint) a tax) 02 a si